OpenStudy (anonymous):

solve the equation 1/x-7+1/x+7=-5/x^2-49

OpenStudy (anonymous):

\[\frac{ 1 }{ x-7 }+\frac{ 1 }{ x+7 }=\frac{ -5 }{ x^2-49 }\] multiply both sides by \[x^2-49 \] and solve for x

OpenStudy (jmark):

1/x-7+1/x+7=-5/x^2-49 take LCM OF LHS equation. (x+7)+ (x-7)/(x-7) (x+7)=-5/x^2-49 by making use of identity (a+b)(a-b) = a2-b2 ((x+7)+ (x-7) )/ (x2-72)=-5/x^2-49 ((x+7)+ (x-7) )/ (x2-49)=-5/x^2-49 multiply both side of the equation by (x^2-49 ) ((x+7)+ (x-7) )=-5 x+7+x-7 = -5 2x = -5 x = -5/2